OpenStudy (anonymous):

If the measure of WZX is 262, what is the measure of XWY?

OpenStudy (anonymous):

@JamesJ

OpenStudy (anonymous):

262/2 ? @JamesJ

OpenStudy (anonymous):

I got 131

OpenStudy (jamesj):

No, that can't be right as the angle is clearly acute, i.e., less than 90 degrees. So what's gone wrong?

OpenStudy (anonymous):

Oh wait do you divide by 360 then? @JamesJ

OpenStudy (jamesj):

Go back one or two questions when we had the other problem like this. Look very carefully at what arc angle we knew and the angle we were calculating. They corresponded somehow. Does that correspondance apply here? If not, how do you get back to that correspondance.

OpenStudy (anonymous):

I think last time we just divided 142 by 2 and got 71..I'm confused on this one @JamesJ

OpenStudy (jamesj):

Yes, to apply that formula what arc length do we need? We need the arc length NOT of WZX but of WX.

OpenStudy (anonymous):

So how do we find that? @JamesJ

OpenStudy (anonymous):

Sorry I'm a bit slow :(

OpenStudy (jamesj):

Weeellll ... what's the total angle of a circle?

OpenStudy (anonymous):

262! :) @JamesJ

OpenStudy (jamesj):

No. 360 degrees. Hence WZX + WX = 360

OpenStudy (jamesj):

and therefore WX = 360 - WZX Make sense?

OpenStudy (anonymous):

Ohhhhh so I do 360 - 262? @JamesJ

OpenStudy (jamesj):

Yes WX = 360 - 262 = 98 Therefore the angle you are asked for in the question is equal to what?

OpenStudy (anonymous):

equal to XWY? @JamesJ

OpenStudy (jamesj):

XWY = WX/2

OpenStudy (anonymous):

49 @JamesJ

OpenStudy (jamesj):

Yes. I'm out of here now.